# Equal Contribution

### v2

@@ 0.A.1 Physical-Aware Losses and Metric
input : Object’s point cloud $\mathcal{P}$, predicted missing part’s mesh $x$, set of position candidate $C_{1}$ determined by grid search, set of small movement $C_{2}$, and loss margin $\beta$.
output : Whether the predicted mesh $x$ is physically plausible
@@ 8            Add $c$ to $C_{0}$;
0.A.2 Experimental Configuration
Pose Proposal Module. The Pose Proposal Module is trained separately for each part category. We train for 100 epochs on a single NVIDIA RTX 4090 with batch size 8. We adopt the Adam optimizer and a $10^{-4}$ learning rate. The architecture is the same as GenPose [78].
3D Vector Quantizer. The 3D-VQVAE model is trained using all part instances from the GAPartNet [17] dataset and is shared across all part categories. We train the model for 150000 iterations on 8 NVIDIA V100 GPUs with batch size 1. We adopt the Adam optimizer and cosine learning rate scheduler with a base learning rate $10^{-5}$. Detailed architecture is shown in Fig. 8.
Part Generation Module. The part generation diffusion module is trained separately for each part category. We train the model for 100000 iterations on 8 NVIDIA V100 GPUs with batch size 4. We adopt the Adam optimizer and cosine learning rate scheduler with a base learning rate $10^{-5}$. Guidance weights are set to $\omega_{1}=3$ and $\omega_{2}=1$ during loss-guided sampling. Fig. 8 illustrates the details.
Part Manipulation Interaction Policy. (1) Slider Drawer: To retrieve items from an open drawer, the gripper moves along the z-axis. To open a drawer, it approaches along the x-axis, typically aiming to grasp a handle on the drawer’s front face. (2) Hinge Door: For hinge doors with front-facing handles, the gripper grabs the handle to open the door. After grasping the handle, it rotates around the predicted hinge axis to complete the opening or closing action. For doors without handles, if the door is ajar, the gripper clamps the outer edge along the y-axis of the bounding box to open it. (3) Hinge Knob: For hinge knobs, the gripper clamps the knob like a round handle and rotates the end-effector to perform the task. (4) Fixed Handle: The interaction policy for linear fixed handles mirrors that of round fixed handles. The gripper’s opening direction should be perpendicular to the linear fixed handle, aligning parallel to the y-axis of the predicted bounding box.
Part Manipulation Interaction Experiments. We employ a heuristic-based interaction policy to open drawers and doors, manipulate handles, and rotate knobs. Once we determine the part pose, our policy immediately provides the corresponding grasping pose. We then use cuRobo [63] to position our gripper at the grasping pose. Following this, our interaction policy and predicted axis guide the design of the end-effector trajectory to match the part’s movement trajectory, interpolating it with a time step of $\frac{1}{200}$. Using Inverse Kinematics and a PID controller, we compute joint poses and maneuver the end-effector along the defined trajectory. Our implementation is independent of ROS, ensuring compatibility with various simulators.
Figure 8: Architecture of our proposed physically plausible generation framework.
Appendix 0.B Additional Experiments
0.B.1 Additional Quantitative Results
Ablation studies on generating dependent parts. Tab. 4 presents ablation studies on dependent parts generation to further validate our design choice to incorporate bounding box conditions into the diffusion model and employ a diffusion-based architecture for pose prediction. Similar to the results for generating self-moving parts shown in Tab. 3, the inclusion of both designs consistently yields superior results, further justifying the effectiveness of our approach in physical-plausible generation.
Table 4: Ablation studies on the pose proposal module and the conditional generation module for dependent parts.
Method Line handle Hinge knob
CD $\downarrow$ F-score $\uparrow$ Phys. (%) CD $\downarrow$ F-score $\uparrow$ Phys. (%)
Regress. pose proposal 0.00252 0.655 74.2 0.00036 0.847 81.1
w/o bbox condition 0.00784 0.572 67.0 0.00033 0.885 80.7
Full version 0.00173 0.744 85.9 0.00024 0.917 84.2
0.B.2 Additional Qualitative Results
Single Part Generation. We present additional single part generation results in Fig. 10 and 11. While the baseline method frequently struggles to generate parts that meet physical constraints, our approach consistently yields high-fidelity part-generation results.
@@ 0.B.2 Additional Qualitative Results
Figure 9: Comparison between whether or not using the bounding box as an additional condition. Incorporating the bounding box beside the point cloud enhances the fit of the generated part to the given object.
Figure 10: Qualitative single part generation results.
Figure 11: Qualitative single part generation results.

### v3

@@ III-B Loss-Guided Physical Constraints
input : Object point cloud $\mathcal{P}$, mesh of the predicted missing part $x$, set of position candidate $C_{1}$ determined by grid search, set of small movement $C_{2}$, and loss margin $\beta$.
output : Whether mesh $x$ is physically plausible
@@ 8            Add $c$ to $C_{0}$;
IV Experiments
This section demonstrates the effectiveness of our method in generating four types of common articulated objects: slider drawers, hinge doors, hinge knobs, and line handles, where the first two allow self-movements and the latter two must be attached to other parts (denoted as dependent parts). We then explore its potential applications in 3D printing, robot manipulation, and sequential part generation tasks.
Data and Implementation. We train our models on the GAPartNet dataset [16], a part-centric object model dataset with part-level annotations and actionability alignment across part classes. Approximately 90% of instances are allocated for training for each part category, with the remainder reserved for evaluation. All part categories utilize the same VQVAE model trained across all part categories at an input SDF resolution of 128, while a separate diffusion model is trained for each category. Both our method and all baselines involve random rotation of input point clouds around the vertical axis during training for fair comparisons. For testing, we sample three different rotation angles for each object instance.
Baselines. Our primary baseline is the Conditional Variational Autoencoder (C-VAE) [54], which similarly utilizes a scale-agnostic SDF representation. To ensure fair comparisons, we scale the generated parts using our predicted pose and then use the same settings to process the results. We also include our method’s performance without the proposed loss-guided sampling to demonstrate its effectiveness.
IV-A Physics-Plausible Generation
Metrics. We introduce a physical plausibility metric to evaluate the viability of generated shapes. As illustrated in Algorithm 1, we first preprocess the meshes, excluding those with multiple large connected parts or those that are not watertight as physically infeasible. For the remaining meshes, a grid search around the part’s ground truth position is conducted to locate positions where both collision loss (Eq. 5) and contact loss (Eq. 6) fall below a specified threshold. A part is deemed physically plausible if its movement along the $+x$ and $-x$ axes (for slider drawers) or along the $-z$ axis (for other part categories) causes these losses to exceed the threshold. Notably, for hinge doors, a shared rotation axis is used during loss-guided sampling, while the ground truth rotation axis is used for evaluation. Besides, we also employ the Chamfer Distance (CD) and F-score [60] (@2%) for evaluation.
Self-moving Part Generation. As shown in Tab. I, our method significantly outperforms the baseline in terms of both generation quality and physical plausibility on generating self-moving parts. The results also highlight the essential role of our proposed physical guidance in the generation process. Notably, as illustrated in Fig. 3, while the baseline method or the absence of physical guidance often leads to collisions or misalignments at the interface between the object and the generated part, our approach consistently delivers parts that seamlessly integrate with the original object, establishing a robust foundation for subsequent dependent part generation.
Dependent Part Generation. As shown in Tab. I, our approach continues to markedly surpass the baseline on dependent parts. Fig. 4 reveals that while the baseline method frequently produces a degenerated distribution (such as generating oval-shaped handles regardless of input or producing broken knobs), our method achieves consistent high-fidelity results that meet the physical requirements.
@@ IV-A Physics-Plausible Generation
Figure 6: Comparison of incorporating the bounding box as condition. Adding the bounding box alongside the point cloud improves the alignment of the generated part with the target object.
IV-B Downstream Applications
3D Printing. To better illustrate the effectiveness and physical plausibility of our model, we 3D-printed the generated meshes. The printed parts not only fit seamlessly into the corresponding articulated objects but also demonstrated highly effective motions. Figure 5(a) shows these 3D-printed shapes, which were scaled down to $\frac{1}{200}$ of their original size, resulting in dimensions of approximately $10\text{cm}\times 10\text{cm}\times 10\text{cm}$.
Simulation for Robot Manipulation. To further validate the physical plausibility and quality of our generated object, we demonstrate that the generated parts can be manipulated using simple heuristics. Following the setup in [16, 15], we load the objects in URDF format into the IsaacGym simulator [38] and apply a heuristic policy based on the estimated pose. Fig. 5(c) shows that our generated shapes exhibit high quality and can be successfully manipulated.
IV-C Ablation Studies and Analysis
This section delves into the effectiveness of different model components through ablation studies.
Bounding Box Condition. We first verify the design of adding the bounding box condition to the diffusion model. As shown by the results in Tab. II and Fig. 6, relying solely on the point cloud condition often yields part generation results that fail to fit the input object, such as drawers colliding or doors not filling the whole empty space. Conversely, incorporating the bounding box as an auxiliary condition significantly improves the overall performance of our model.
Pose Proposal Model. We then verify the effectiveness of adopting the diffusion-based pose proposal model. As also shown in Tab. II, leveraging the diffusion-based structure produces object bounding boxes of higher accuracy, thus benefiting the overall part generation quality.
V Conclusions
In this paper, we propose a diffusion-based framework that integrates geometric conditioning and physical constraints, ensuring high accuracy and fidelity of the generated parts in shape and function under real-world interactions. Our results demonstrate substantial improvements in shape quality and physical plausibility over existing methods. These enhancements significantly benefit downstream applications such as 3D printing and robot manipulation, where precision is crucial for the tasks’ success.